What fraction of a clockwise revolution does the hour hand of a clock turn through, when it goes from
(a) 3 to 9
(b) 4 to 7
(c) 7 to 10
(d) 12 to 9
(e) 1 to 10
(f) 6 to 3

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) 3 to 9 9 – 3 = 6 ÷ 12 = 12 of a revolution (b) 4 to 7 7 – 4 = 3 ÷ 12 = 14 of a revolution (c) 7 to 10 10 – 7 = 3 ÷ 12 = 14 of a revolution (d) 12 to 9 i.e., 0 to 9 9 – 0 = 9 ÷ 12 = 34 of a revolution (e) 1 to 10 10 – 1 = 9 ÷ 12 = 34 of a revolution (f) 6 to 3 i.e., 6 to 12 and then 12 to 3 6 to 12 = 12 – 6 = 6 and 12 to 3 = 0 to 3 = 3 – 0 = 3 6 + 3 = 9 ÷ 12 = 34 of a revolution